OpenStudy (mackenzie2013):

The GCF of 14 and a number is 7. Which of the following could be the number? 28 42 49 56

OpenStudy (anonymous):

49: 28/2=14 42/3=14 56/4=14 49=7^2

OpenStudy (mackenzie2013):

Thank you :)

OpenStudy (skullpatrol):

Did you understand anything @Mackenzie2013 ?

OpenStudy (mackenzie2013):

I did , Thank you!